1/1

William Wordsworth,  William Wordsworth,
The Daffodils  Dafodiloj

tradukita de Krys Ungar

I wander’d lonely as a cloud That floats on high o’er vales and hills, When all at once I saw a crowd, A host of golden daffodils, Beside the lake, beneath the trees Fluttering and dancing in the breeze.

Mi vagis sola kiel nub’ Altflosas super la kampar’ Kaj tiam tuj aperis trup’ De dafodiloj; ora ar’ Apud la lag’, sub arba tend’ Flirtis kaj dancis en la vent’.

Continuous as the stars that shine And twinkle on the milky way, They stretch’d in never-ending line Along the margin of a bay: Ten thousand saw I at a glance Tossing their heads in sprightly dance.

Senbreĉe kiel astra bril’ Sur Lakta Voj’ flagretas, jen Sterniĝis ili en defil’ Senfine laŭ la laghaven’. Rigardon kaptis la balanc’ De dekmil kapoj en la danc’.

The waves beside them danced, but they Out-did the sparkling waves in glee: - A poet could not but be gay In such a jocund company: I gazed - and gazed - but little thought What wealth the show to me had brought.

Apudaj ondoj dancis, sed Triumfon gajnis floroj ĉi, Devigis gajon de poet’ Plej joviala kompani’: Mi rigardadis sen atent’, Ke min riĉigos la prezent’.

For oft, when on my couch I lie In vacant or in pensive mood, They flash upon that inward eye Which is the bliss of solitude; And then my heart with pleasure fills And dances with the daffodils.

Ĉar ofte, kiam kuŝas mi

Penseme, aŭ en menslangvor’, Ekfulmas ili tra l’ konsci’ - La beno de solema hor’; La koro pro pleniga ĝu’ Kun dafodiloj dancas plu.

Verkinto de tiu ĉi Angla poemo estas William Wordsworth (∗1770 – †1850).

Arg-39-73 (2003-10-17 22:02:21)

Origina anglalingva teksto de WILLIAM WORDSWORTH (1770 - 1850)

Traduko de la Angla poemo “The Daffodils” de William Wordsworth (∗1770 – †1850) en Esperanton de Krys Ungar (Krys Bottrill).

Arg-39-632 (2009-09-20 15:01:36)

http://www.poezio.net/version?poem-id=39&version-id=73,632